OpenStudy (anonymous):

9 squared divided by 3 cubed. Give expression in symbols and evaluate it? What is a exponential mot action with positive exponents?

OpenStudy (anonymous):

\[9^{2}/3^{3}=9*9/3*3*3=3*3*3*3/3*3*3=3\]

OpenStudy (anonymous):

I'm not understanding what is being said? So what would the given expression in symbols be?

OpenStudy (anonymous):

\[9^{2}/3^{3}\] this is the expretion in symbols

OpenStudy (anonymous):

Okay would I put the divided sign

OpenStudy (anonymous):

@myko has answered the question, and shown the symbolic notation with one minor change if I may: \[\frac{9^{2}}{3^{3}}=\frac{9*9}{3*3*3}=\frac{3*3*3*3}{3*3*3}=3\] 3*3*3*3/3*3*3=3*3*3*1*3*3=243 (3*3*3*3)/(3*3*3) =3 Parenthesis please ladies & gents! :-D

OpenStudy (anonymous):

or use fraction: \[9^{2}\over 3^{3}\]

OpenStudy (anonymous):

(9^2/3^3)